# Active Advantage-Aligned Online Reinforcement Learning with Offline Data

### v3

@@ Introduction
Density ratio. We evaluate the onlineness through the use of a density ratio
@@
for a given transition, where $d^{\text{on}}\left({s,a}\right)$ denotes the state-action distribution of online samples in the online buffer $\mathcal{R}^{\text{on}}$ and the $d^{\text{off}}\left({s,a}\right)$ denotes the offline samples in the offline buffer $\mathcal{R}^{\text{off}}$. By identifying a transition with a high density ratio ${w}\left({s,a}\right)$, we can effectively select a near-on-policy sample $\left({s,a,s^{\prime}}\right)$ from the offline dataset $\mathcal{B}^{\text{off}}$. Consider the much larger volume of offline data compared to online data, this would greatly improve the amount of transition and diversity of coverage for policy improvement in each step.
Estimating the likelihoods $d^{\text{off}}\left({s,a}\right)$ and $d^{\text{on}}\left({s,a}\right)$ poses a challenge, as they could represent stationary distributions from mixture of complex policy. To address this issue, we employ a method studied by Lee et al. (2022); Sinha et al. (2022) for density ratio estimation that does not rely on likelihoods. This method approximates ${w}\left({s,a}\right)$ by training a neural network ${w}_{\psi_{i}}\left({s,a}\right)$, which is parameterized by $\psi_{i},i\in[N_{e}]$, where $N_{e}$ is the number of density networks in the ensemble. The training exclusively uses samples from $\mathcal{B}^{\text{off}}$ and $\mathcal{B}^{\text{on}}$. We use variational representation of f-divergences (Nguyen, Wainwright, and Jordan 2007). Consider $P$ and $Q$ as probability measures on a measurable space $\mathcal{X}$, with $P$ being absolutely continuous $w.r.t$ $Q$. We define the function $f\left({y}\right):=y\log\frac{2y}{y+1}+\log\frac{2}{y+1}$. The Jensen-Shannon (JS) divergence is then defined as $D_{JS}\left({P||Q}\right)=\int_{\mathcal{X}}f\left({{dP\left({x}\right)}/{dQ\left({x}\right)}}\right)dQ\left({x}\right)$. Then we use a parametric based model ${w}_{\psi}\left({x}\right)$ to represent density ratio $\frac{dP}{dQ}$ and estimated the density ratio by maximizing the lower bound of $D_{JS}\left({P||Q}\right)$:
$\mathcal{L}^{\text{DR}}\left({\psi}\right)=\mathbb{E}_{x\sim P}\!\left[f^{\prime}\left({{w}_{\psi}\left({x}\right)}\right)\right]-\mathbb{E}_{x\sim Q}\!\left[f^{*}\left({f^{\prime}\left({{w}_{\psi}\left({x}\right)}\right)}\right)\right],$
@@
4:Initialize buffer $\mathcal{D}$ with offline data, online replay buffer $\mathcal{R}\leftarrow\emptyset$
5:while True do
6:  Receive initial observation state $s_{0}$
7:  for $t=0,\ldots,T$ do
8:   Take action $a_{t}\sim\pi_{\phi}\left({\cdot|s_{t}}\right)$, update buffer $\mathcal{R}\leftarrow\mathcal{R}\cup\left\{\left({s_{t},a_{t},r_{t},s_{t+1}}\right)\right\}$.
9:   Randomly sample a subset of size $\frac{N}{2}$ from online buffer $\mathcal{R}$ and size $\frac{N}{2}$ from offline buffer $\mathcal{D}$ to form a learning dataset $\mathcal{R}_{N}$
10:   Update density ensemble using $\mathcal{R}_{N}$
11:   Calculate priority $P_{\mathcal{R}}$ of $\mathcal{R}_{N}$ via (4)
12:   for $g=1,\ldots,G$ do
13:     Sample batch $b_{N}$ of size NN according to $P_{\mathcal{R}}$ from $\mathcal{R}_{N}$
14:     Sample set $\mathcal{Z}$ of $Z$ indices from $\left\{1,\ldots,E\right\}$
15:     With $b_{N}$, set $y=r+\gamma\Big{(}\min_{i\in\mathcal{Z}}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)+\Big{.}$ $\Big{.}\alpha\log\pi_{\phi}\left({a^{\prime}|s^{\prime}}\right)\Big{)},a^{\prime}\sim\pi_{\phi}\left({\cdot|s^{\prime}}\right)$
16:     for $i=1,\ldots,E$ do
17:      Calculate importance weight $u_{i}$ via (3).
18:      Update $\theta_{i}$ minimizing loss: $\ell=\sum_{i}u_{i}\cdot\left({y-Q_{\theta_{i}}\left({s,a}\right)}\right)^{2}$
19:     Update target networks: $\theta^{\prime}_{i}\leftarrow\rho\theta^{\prime}_{i}+\left({1-\rho}\right)\theta_{i}$
20:   With $b_{N}$, update $\phi$ maximizing objective:
21:     $\frac{1}{E}\sum_{i=1}^{E}Q_{\theta_{i}}\left({s,a}\right)-\alpha\log\pi_{\phi}\left({a|s}\right),$ where $a\sim\pi_{\phi}\left({\cdot|s}\right)$, $\left({s,a}\right)\sim b_{N}$.
where ${w}_{\psi}\left({x}\right)\geq 0$ is represented by a neural network, with parameters ensuring that the outputs remain non-negative through the use of activation function. Additionally, $f^{*}$ represents convex conjugate and we sampled from $\mathcal{B}^{\text{on}}$ for $x\sim P$ and from $\mathcal{B}^{\text{off}}$ for $x\sim Q$.
@@
Relying solely on the density ratio is insufficient; even if a transition appears to be relevant in the online context, it may still fail to contribute meaningfully to policy improvement. For instance, consider a transition $\left({s,a,s^{\prime}}\right)$. If the policy has previously encountered this state and taken the same action, or if the action performed in this state could potentially lead to a negative reward, such a transition would not that helpful in contributing to policy improvement, regardless of how closely it aligns with on-policy data.
To address this, we incorporate an estimate of the advantage value $A(s,a)$ (Eq. (2)) into our sampling strategy. Specifically, we integrate a non-negative exponential advantage term, $\exp\left({\mathcal{\xi}\cdot A\left({s,a}\right)}\right)$, into the priority calculation. This term ensures that transitions are selected not only based on relevance but also on their contribution to policy improvement. The higher the advantage value, the greater the transition’s impact on learning, making our sampling mechanism both adaptive and optimization-aware.
@@
For transitions from the offline dataset, we prioritize samples based on both the estimated density ratio and advantage value, retrieving near-on-policy samples that also provide policy improvement benefits. Since the data source is known, we set the density ratio to 1 for transitions from the online dataset and prioritize them purely based on advantage values under the current policy. Additionally, there may be uncertainty and significant variance in estimating the advantage value and density ratio. To address this, we adopt LCB as a conservative estimate. Thus, we define the priority function for sampling as:
$\mathbb{I}^{\text{off}}\underline{{w}}\left({s_{i},a_{i}}\right)\cdot\exp\left({\mathcal{\xi}\cdot\underline{A}\left({s_{i},a_{i}}\right)}\right)+\mathbb{I}^{\text{on}}\exp\left({\mathcal{\xi}\cdot\underline{A}\left({s_{i},a_{i}}\right)}\right).$
@@
Note that this advantage-aligned sampling strategy is not a heuristic-based approach but is theoretically derived in the performance difference lemma (Kakade and Langford 2002), providing insights into its effectiveness and superiority over the random sampling approach (see Section Theoretical Analysis ).
@@
The active sampling process in our algorithm is highlighted in blue in Algorithm 1, while our approach to addressing sampling bias is highlighted in red.
Theoretical Analysis
@@
In this section, we derive the priority term theoretically from the performance difference lemma (Kakade and Langford 2002) and show that our active advantage-aligned sampling strategy leads to improved policy performance. Furthermore, we establish a theoretical lower bound on the performance improvement gap under our sampling scheme.
@@
The proof is provided in the Appendix. We note that the coefficient $R^{t}(s,a;\xi)$ is not necessarily the tightest possible bound, since it is based on the supremum norm and therefore can be dominated by a single $(s,a)$ pair. A sharper result could be obtained by measuring distribution shift in the $\ell_{2}$ norm (or some other weaker norm). We nevertheless adopt the simpler supremum-norm bound here for clarity and to highlight the core intuition behind why advantage reweighting yields improvement, as will be detailed in the following.
a halfcheetah-random
b ant-random
c walker2d-medium
d antmaze-medium
Figure 2: Main results. A comparison between $A^{3}$RL, the state-of-the-art baseline RLPD and SAC with offline data (SACfD) on various D4RL benchmark tasks. (a-c): dense reward, (d): sparse reward. Shaded areas represent one standard deviation based on ten seeds.
Comparison to random sampling.
The fundamental concept behind proving that our sampling technique surpasses random sampling and contributes to positive policy improvement involves initially applying the performance difference lemma. This approach yields the performance differential term $J\left({\pi^{t+1}}\right)-J\left({\pi^{t}}\right)$ between the updated policy and the current policy. Our goal is to demonstrate that this term is non-negative under our sampling priority. To do this, we prove that by a shift of distribution, this term is no less than the gap
@@
Why does advantage weighting help?
We show that under certain conditions, the ratio $R^{t}(s,a)$ can decrease for increased value of $\xi$. Since $\xi$ does not influcence the ration between the state distribution, let us just consider the bandit case with ratio
@@
Experiments
a advantage term
b density term
c purely online setting
d TD-error priority
Figure 3: Ablation Studies: Results of ablation studies on the halfcheetah-random environment.
Figure 4: halfcheetah-expert
Environments.
We evaluate $A^{3}$RL on both dense and sparse reward tasks from the D4RL benchmark (Fu et al. 2020). These include halfcheetah, walker2d, and ant, which are dense reward locomotion tasks, and antmaze, which involves sparse rewards. Each environment offers offline datasets composed of trajectories ranging from completely random to expert. We defer additional details on the environment to the Appendix.
Setup.
We employ the basic setup of the SAC networks as recommended by (Ball et al. 2023), i.e., with an ensemble of size 10 each for critic networks and target critic networks, as well as entropy regularization. A significant difference is that the MLP underlying these networks only has 2 layers of size 256 each, as we desired to see if the agent is able to learn with less complexity.
Baseline Methods.
For our main results, we compare $A^{3}$RL with two baselines: (1) RLPD (Ball et al. 2023), regarded as the SOTA baseline for addressing online RL with offline datasets, also attains state-of-the-art performance in this problem set, (2) SAC with offline data (SACfD), a canonical off-policy approach using offline data, as also studied in (Nair et al. 2020) and (Vecerik et al. 2017). In the ablation studies, we evaluate $A^{3}$RL against five additional representative baselines: (3) Off2On (Lee et al. 2022), an offline-to-online RL method; (4) a variant of $A^{3}$RL using advantage estimation only; (5) an online version of $A^{3}$RL that excludes offline data; (6) SAC in an online setting without offline data; (7) TD (Temporal Difference) with a PER (Schaul et al. 2015) sampling strategy; and (8) TD+Density, which combines PER with a density ratio sampling strategy.
Main results
Fig. 2 presents a comparative analysis of $A^{3}$RL’s performance against the baseline SACfD and the current state-of-the-art method, RLPD. The results demonstrate that $A^{3}$RL consistently outperforms the baseline across the evaluated domains. This performance advantage can be attributed to a fundamental difference in sampling strategy: while RLPD relies on symmetric random sampling, $A^{3}$RL employs an active sampling approach based on advantage alignment.
Unlike RLPD, which treats all transitions uniformly, $A^{3}$RL dynamically reevaluates the relevance and on-policyness of each transition as the policy evolves, continuously adjusting its sampling priority to align with the current learning needs. This targeted sampling ensures that the most beneficial transitions are prioritized, directly contributing to faster and more effective policy improvement.
In scenarios involving nearly random offline datasets  Fig.2a,2b, datasets containing trajectories from a poorly performing policy, or even medium datasets Fig.2d,2c, useful transitions are often sparse and scattered. Random sampling, as used by RLPD, is likely to miss these valuable data points, leading to suboptimal performance. In contrast, $A^{3}$RL ’s active sampling strategy effectively identifies and emphasizes these critical transitions, resulting in substantial policy enhancements, as clearly illustrated in Fig. 2.
In expert environments Fig. 4, RLPD performs on par with $A^{3}$RL. This superior performance can be attributed to the higher quality of transitions present in medium and expert datasets, compared to random datasets. Consequently, even with a random sampling strategy, RLPD is still likely to encounter useful transitions. However, most offline datasets are provided in a black-box format, where the specifics of the data are unknown. Despite this uncertainty, $A^{3}$RL achieves performance that is at least on par with RLPD, demonstrating robustness to varying data quality in these black-box conditions.
Ablation studies
Ablation on advantage term.
Fig. 3a illustrates the comparison between the performance of $A^{3}$RL using advantage-aligned sampling priority and Off2On utilizing solely density ratio (${\sigma}=\mathbb{I}^{\text{off}}{w}\left({s_{i},a_{i}}\right)+\mathbb{I}^{\text{on}}$), a modified version of balanced experience replay (Lee et al. 2022). The results show that $A^{3}$RL with the advantage term surpasses its counterpart that only considers online-ness in prioritizing samples in the halfcheetah-random environments. This superiority is attributed to the advantage term, which effectively screens out transitions that are either non-informative or harmful. For example, even if a transition indicates online-ness, it may not provide new information if the policy has already mastered the associated action for that state. By integrating the advantage term, such repetitive transitions are excluded, as the advantage value tends to zero for well-understood transitions.
Ablation on density term.

### v4

@@ 4 Algorithm
To evaluate the onlineness/on-policyness of an offline transition, we estimate the density ratio
@@
where $d^{\text{on}}\left({s,a}\right)$ denotes the state-action density of online samples in the online buffer $\mathcal{R}$ and the $d^{\text{off}}\left({s,a}\right)$ denotes that of the offline dataset $\mathcal{D}$. Then, by identifying a transition with a high density ratio ${w}\left({s,a}\right)$, we can effectively select a near-on-policy transition $\left({s,a,s^{\prime}}\right)$ from the offline dataset $\mathcal{D}$. This directly leverages the abundance of the offline data compared to online data, greatly improving the amount of transitions and diversity of coverage for policy improvement in each step.
Estimating the likelihoods $d^{\text{off}}\left({s,a}\right)$ and $d^{\text{on}}\left({s,a}\right)$ poses a challenge, as they could represent stationary distributions from mixture of complex policy. To address this issue, we employ a method studied by Lee et al. [30] and Sinha et al. [59] for density ratio estimation that does not rely on likelihoods, namely by approximating ${w}\left({s,a}\right)$ with a neural network ${w}_{\psi}\left({s,a}\right)$. To this end, we use variational representation of f-divergences [45] in the following way. Let $P$ and $Q$ be probability measures on a measurable space $\mathcal{X}$, with $P$ being absolutely continuous w.r.t. $Q$. Let $f\left({y}\right):=y\log\frac{2y}{y+1}+\log\frac{2}{y+1}$, then the Jensen-Shannon (JS) divergence between $P$ and $Q$ can be written as $D_{\textrm{JS}}\left({P||Q}\right)=\int_{\mathcal{X}}f\left({(dP/dQ)(x)}\right)dQ\left({x}\right)$. It then has the variational lower bound
$\mathcal{L}^{\text{DR}}\left({\psi}\right)=\mathbb{E}_{x\sim P}\!\left[f^{\prime}\left({{w}_{\psi}\left({x}\right)}\right)\right]-\mathbb{E}_{x\sim Q}\!\left[f^{*}\left({f^{\prime}\left({{w}_{\psi}\left({x}\right)}\right)}\right)\right],$ (5)
where $f^{*}$ denotes the convex conjugate of $f$, that is maximized at $w_{\psi}(x)=(dP/dQ)(x)$. This enables us to use a parametric model ${w}_{\psi}\left({x}\right)$ to approximate the density ratio $\frac{dP}{dQ}$ via maximizing this $\mathcal{L}^{DR}$ lower bound of $D_{\textrm{JS}}\left({P||Q}\right)$.
@@
1:Select ensemble size $E$, UTD ratio $G$, discount $\gamma$, temperature $\alpha$, critic polyak weight $\rho$, batch size NN.
2:Randomly initialize critic $\left\{\theta_{i},i\in\left[E\right]\right\}$ and actor $\phi$ parameters. Set targets $\theta^{\prime}_{i}=\theta_{i}$.
3:Initialize offline priority buffer $\mathcal{B}^{\text{off}}$ with offline dataset $\mathcal{D}$, online priority buffer $\mathcal{B}^{\text{on}}=\mathcal{R}\leftarrow\emptyset$ .
4:for each environment step $t$ do
5:  Take action $a_{t}\sim\pi_{\phi}\left({\cdot|s_{t}}\right)$, update online buffer $\mathcal{B}^{\text{on}}\leftarrow\mathcal{B}^{\text{on}}\cup\left\{\left({s_{t},a_{t},r_{t},s_{t+1}}\right)\right\}$.
6:  for $g=1,\ldots,G$ do
7:   Uniformly sample $\frac{N}{2}$ transitions from $\mathcal{B}^{\text{off}}$ and similarly $\frac{N}{2}$ from $\mathcal{B}^{\text{on}}$. Update $w_{\psi}$ to maximize (5).
8:   Sample with priority $\frac{N}{2}$ transitions from $\mathcal{B}^{\text{off}}$ and similarly $\frac{N}{2}$ from $\mathcal{B}^{on}$ to form batch $b$ of size NN.
9:
10:   With $b$, randomly subsample $\mathcal{Z}\subset[E]$ with $\left|\mathcal{Z}\right|=2$, set
$y=r+\gamma\Big(\min_{i\in\mathcal{Z}}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)+\alpha\log\pi_{\phi}\left({a^{\prime}|s^{\prime}}\right)\Big),\quad a^{\prime}\sim\pi_{\phi}\left({\cdot|s^{\prime}}\right).$
11:   Within each buffer, calculate importance weights $u$ via (3).
12:   Scale weights so that sum of offline weights = sum of online weights = $1/2$.
13:   Perform SAC critic updates with target $y$ and target critic polyak updates.
14:  Perform SAC actor updates and update priorities for $b$ according to (6).
In practice, since ${w}_{\psi}\left({x}\right)$ approximates $\frac{dP}{dQ}\geq 0$, it has to be constrained to be non-negative (through the use of appropriate activation functions). Furthermore, since we’re estimating $d^{\text{on}}/{d^{\text{off}}}$, for $\mathcal{L}^{\text{DR}}$ we sample from $\mathcal{R}$ for $x\sim P$ and from $\mathcal{D}$ for $x\sim Q$. To get more stable density ratio estimates $\underline{w}$, similar to Lee et al. [30], we self-normalize estimated density ratio $w$ within $\mathcal{D}$ with a temperature hyperparameter $\zeta$:
$\underline{w}_{\psi}(x)={w_{\psi}(x)^{\zeta}}/{\mathbb{E}_{y\sim\mathcal{D}}\left[w_{\psi}(y)^{\zeta}\right]}.$
Advantage term.
For the advantage term, to enhance robustness, we use the pessimistic CDQ $Q$ estimation through random subsampling of an ensemble $\left\{Q_{\theta_{i}},Q_{\theta^{\prime}_{i}}:i\in[E]\right\}$ of $Q$ and target $Q$ networks where $E$ is the ensemble size, similar to Ball et al. [2], while also incorporating uncertainty estimation for the advantage in the following way. Since $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)=A^{\pi}(s,a)=Q^{\pi}(s,a)-\mathbb{E}_{a^{\prime}\sim\pi(s)}\left[Q^{\pi}(s,a^{\prime})\right]$, we can get $E$ noisy estimates of the advantage by simply taking $A_{i}(s,a)=Q_{\theta_{i}}(s,a)-Q_{\theta_{i}}(s,a^{\prime})$ where $a^{\prime}\sim\pi$. The estimated advantage we then use is the advantage lower confidence bound (LCB) $\underline{A}(s,a)$, calculated as
$\underline{A}(s,a)=\widehat{A}(s,a)-\beta\widehat{\sigma}(s,a),$
where $\widehat{A}(s,a)$ and $\widehat{\sigma}(s,a)$ are respectively the sample mean and standard deviation of $\left\{A_{i}(s,a):i\in[E]\right\}$, and $\beta$ is a hyperparameter confidence level derived from the Student’s $t$-distribution. The immediate upside of this calculation is that quantities like $Q_{\theta_{i}}(s,a)$ and $Q_{\theta_{i}}(s,a^{\prime})$ are already computed during regular SAC updates, making this calculation computationally inexpensive.
@@
Here, we make the argument that relying solely on either of the two aforementioned terms would yield inevitable insufficiencies. Relying solely on the density ratio is insufficient; even if a transition appears to be relevant in the online context, it may still fail to contribute meaningfully to policy improvement, e.g., it might simply represent a "bad" transition in the value sense. Consider a transition $\left({s,a,s^{\prime}}\right)$. If the policy has previously encountered this state and taken the same action, or if the action performed in this state could potentially lead to a negative reward, such a transition would not be that helpful in contributing to policy improvement, regardless of how closely it aligns with on-policy data. On the other hand, as lengthily discussed in prior works (e.g. Lee et al. [30]), not addressing the offline-online distributional shift could also yield catastrophic results.
Therefore, we incorporate both estimated advantage and density ratio into our sampling strategy, ensuring that transitions are selected not only based on relevance, but also on their contribution to policy improvement. Concretely, this is done via setting a transition’s priority to
$p(s,a)=\begin{cases}\underline{{w}}(s,a)\exp(\xi\underline{A}(s,a))&\text{ if }(s,a)\in\mathcal{D}\\ \exp(\xi\underline{A}(s,a))&\text{ if }(s,a)\in\mathcal{R}\end{cases},$
or more succinctly,
$p(s,a)=(\mathbb{I}^{\text{off}}\underline{{w}}(s,a)+\mathbb{I}^{\text{on}})\exp(\xi\underline{A}(s,a)),$
where $\mathbb{I}^{\text{off}}$ and $\mathbb{I}^{\text{on}}$ are indicator functions of the offline dataset and online buffer respectively.
First, we use the estimated advantage through the exponential term $\exp(\xi\underline{A}(s,a))$. The natural consequence is that the more advantage a transition has (and thus the greater its impact on learning is), the more priority it is assigned, making our sampling mechanism both adaptive and optimization-aware. Since there may be uncertainty and significant variance in estimating the advantage, we adopt the LCB [36] a conservative estimate, as previously detailed.
@@
Second, we assign an additional weight of $\underline{w}(s,a)$ for transitions from the offline dataset. In other words, we prioritize offline samples based on both its "onlineness" and advantage value to retrieve near-on-policy samples that also provide policy improvement benefits. For the online transitions, they are themselves samples from the online data distribution, and thus we prioritize them purely based on advantage values under the current policy.
Third, we found that the implementation of this prioritized sampling scheme is most straightforward when done as a variation of RLPD [2], where we similarly give equal total weights to the offline dataset and the online replay buffer, but distribute them non-uniformly. This reaps the benefits of RLPD over previous offline-to-online RL methods, namely avoiding the phase separation between offline pretraining and online finetuning, and directly utilizing the offline dataset in conjunction with online interactions instead.
@@
Last but not least, this advantage-aligned sampling strategy is not a heuristic-based approach. It is theoretically derived from the performance difference lemma [23], which provides insights into its effectiveness and superiority over the random sampling approach. For this, see section §5.
In summary, our approach broadens the distribution of samples used during online updates, centering around on-policy examples, thereby facilitating immediate value. The active advantage-aligned priority is as follows:
$p(s,a)$ $=(\mathbb{I}^{\text{off}}\underline{{w}}(s,a)+\mathbb{I}^{\text{on}})\exp(\xi\underline{A}(s,a)),$ (6a)
$\underline{A}(s,a)$ $=\widehat{A}(s,a)-\beta\widehat{\sigma}(s,a),$ (6b)
$\underline{w}_{\psi}(s,a)$ $=\frac{w_{\psi}(s,a)^{\zeta}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\mathcal{D}}\left[w_{\psi}(s^{\prime},a^{\prime})^{\zeta}\right]},$ (6c)
with hyperparameter $\beta$ associated with the LCB advantage estimated, $\zeta$ as the density self-normalization temperature and $\xi$ as the advantage temperature. Note that the standard PER temperature can be absorbed into these hyperparameters, and thus there is no need to define a separate one.
@@
This active sampling process in our algorithm is highlighted in blue in Algorithm 1, while our approach to addressing sampling bias is highlighted in red. We defer the details of the algorithm to Appendix 2.
5 Theoretical Analysis
@@
In this section, we derive the priority term theoretically from the performance difference lemma [23] and show that our active advantage-aligned sampling strategy leads to improved policy performance. Furthermore, we establish a theoretical lower bound on the performance improvement gap under our sampling scheme.
@@
The proof is provided in the Appendix A. We note that the coefficient $R^{t}(s,a;\xi)$ is not a tight bound, since it is based on the supremum norm and therefore can be dominated by a single $(s,a)$ pair. A sharper result could be obtained by measuring distribution shift in the $\ell_{2}$ norm (or some other weaker norm). We nevertheless adopt the simpler supremum‐norm bound here for clarity and to highlight the core intuition behind why advantage reweighting yields improvement, as will be detailed in the following.
a door-expert
b hammer-cloned
c pen-cloned
d pen-human
e relocate-expert
f halfcheetah-medium
g humanoid-medium
h humanoidstandup-medium
Figure 2: Main results. A comparison between $A^{3}$RL in blue, the SOTA baseline RLPD [2] in red, PEX [71] in green and BOORL [19] in orange on various D4RL benchmark tasks. $A^{3}$RL scores the best in all benchmarks, and the gap is especially large for D4RL Adroit tasks (door, hammer, pen, relocate), which are harder due to their larger action dimensionality. Both PEX and BOORL require an offline pretraining process of 1M gradient steps each, and only the online finetuning phase with an initial pretrained jumpstart is shown here. In this view, $A^{3}$RL is much more computationally efficient in achieving the same level of performance.
Comparison to random sampling. The fundamental concept behind proving that our sampling technique surpasses random sampling and contributes to positive policy improvement involves initially applying the performance difference lemma. This approach yields the performance differential term $J\left({\pi^{t+1}}\right)-J\left({\pi^{t}}\right)$ between the updated policy and the current policy. Our goal is to demonstrate that this term is non-negative under our sampling priority. To do this, we prove that by a shift of distribution, this term is no less than the gap
@@
Why does weighting by advantage help? We show that under certain conditions, the ratio $R^{t}(s,a)$ can decrease for increased value of $\xi$. Since $\xi$ does not influence the ratio between the state distribution, let us just consider the bandit case with ratio
@@
6 Experiments
Environments. We evaluate $A^{3}$RL on tasks from the D4RL benchmark [12] with Minari offline datasets [69]. Each environment has offline datasets composed of trajectories of a wide range of competence, ranging from simple to to expert on some, and cloned to expert on others. We defer additional details on the environments to the Appendix F.
Setup. We employ the basic setup of the SAC networks as recommended by [2]. In particular, for critic networks and target critic networks, we have an ensemble of size $E=10$ each, as well as entropy regularization. We use simpler multi-layer perceptrons of 2 layers of size 256 each to see if the agent is able to learn with less model capacity. Runs are compared over 250K online environment steps, with checkpoint evaluation at every 5K steps. Shaded areas represent one standard error of the mean based on ten seeds.
Baseline methods. For our main results, we compare our $A^{3}$RL algorithm against three baselines: (1) RLPD [2], regarded as the SOTA baseline for addressing online RL with offline datasets, specifically having attained SOTA performance in these tasks; (2) PEX [71]; and (3) BOORL [19]. Both PEX and BOORL require an offline pretraining process, and thus only results from their online finetuning phase with an initial jumpstart are shown in our plots. For the training of $A^{3}$RL, we found it beneficial to simply run RLPD for an initial phase of $1/4$ the total number of environment steps and the full $A^{3}$RL for the remaining $3/4$. This enables the $Q$ and density networks to be stabilized for more accurate density and advantage estimations in the absence of an explicit offline pretraining phase.
We then conduct several ablation studies to investigate the sensitivity and effect of each part of our algorithm: (4) ablation on the density term $(\zeta=0)$ yields a variant of $A^{3}$RL that only uses advantage estimations; (5) ablation on the advantage term $(\xi=0)$ yields a density-only-based implementation that is similar to Off2On [30], an offline-to-online RL method, which also does density-based prioritized sampling but requires an offline pretraining phase; (6) ablation on the use of the LCB over the ensemble mean $(\beta=0)$ yields a non-confidence-aware variant of $A^{3}$RL. We also experiment with (7) removing the offline dataset altogether to yield an online version of $A^{3}$RL, which naturally no longer involves the density term; and (8) comparing our prioritization scheme against the commonly used PER scheme. Please refer to Appendix F for our full list of runs on more diverse environments and datasets.
6.1 Main Results
a Density, $\zeta=0$
b Advantage, $\xi=0$
c LCB, $\beta=0$
d Online
e TD
f Offline entropy
g door-cloned
h door-human
Figure 3: Ablation studies and effects of $A^{3}$RL. The unablated version of $A^{3}$RL is in blue throughout. (a)(b)(c) Representative ablation studies of the density term, the advantage term and the LCB estimation. (d) Comparison of the online variant of $A^{3}$RL against purely using SAC online. (e) Comparison of $A^{3}$RL against using TD as the priority term. (f) The typical evolution of the entropy of the prioritized offline buffer. (g)(h) Comparison of $A^{3}$RL against RLPD, PEX, BOORL in the lower-quality/data-starved regime.
Fig. 2 presents a comparative evaluation of $A^{3}$RL against the baselines PEX, BOORL, and the current SOTA method RLPD. The results illustrate that $A^{3}$RL consistently outperforms these methods across all evaluated domains, achieving superior performance, robustness, and training stability.
As consistent in all subplots in Fig. 2, the performance of $A^{3}$RL is essentially identical to that of RLPD in the first $1/4$ of steps. Of course, this is expected because we train $A^{3}$RL with an initial phase of RLPD. However, after this point, their performances quickly diverge. The active sampling scheme of $A^{3}$RL enables it to stay robust in harder D4RL Adroit tasks, such as door-expert and relocate-expert, while RLPD is less sample-efficient and effective. This performance advantage can be attributed to the fundamental difference in the two sampling strategies: while RLPD relies on uniform random sampling, $A^{3}$RL employs an active sampling approach.
Meanwhile, although the other two baselines, PEX and BOORL, promised to address the issues related to catastrophic forgetting and offline-online distribution shifts that come hand-in-hand with an explicit offline pretraining phase, these guarantees are not fulfilled. Oftentimes, the offline pretrained models do not offer a competitive initial jumpstart in these harder tasks (e.g., Fig. 2a, Fig. 2b), which defeat the goals of an explicit, computationally expensive offline pretraining phase. Furthermore, even when they do (e.g., Fig. 2e, Fig. 2f, Fig. 2g), the effects of both catastrophic forgetting and distributional shift are ever present. It is in this light that the benefits of an “end-to-end” approach like $A^{3}$RL and RLPD are clear over computationally expensive and non-robust methods with an offline pretraining phase, at least when we are concerned with sufficiently many online environment steps.
6.2 Ablation Studies and Effects of $A^{3}$RL